# Uncertainty-Aware Probabilistic Graph Neural Networks for Road-Level Traffic Accident Prediction

### v3

@@ 4.1.1 Statistic Analysis
(b) The distribution of accident risk scores frequency in the three boroughs
Figure 2: Traffic accidents are depicted for the case studies.
The three high-accident frequency boroughs of Greater London, UK are illustrated in Figure 2. These boroughs each averaged 3 to 5 daily accidents throughout 2019, as detailed in Table 3. Despite these figures, a significant number of roads in these boroughs report no accidents, leading to a pronounced zero-inflation problem. This situation is compounded by the distribution of accident risk values (Figure 2), which shows a zero-inflated problem along with a pronounced long-tailed effect -where a few high-risk accident occurrences contrast with many low-risk ones. Such data characteristics underscore the suitability of our proposed model for this dataset, effectively capturing both the prevalent zero rates and the diverse spread of accident risk levels.
Table 3: The statistics of Westminster, Lambeth, and Tower Hamlets datasets.
Dataset  Roads/Nodes   Edges   Accident Counts   Zero-Inflated
@@
4.1.2 Lower-Dimensional Representation for Traffic Accident
To interpret the representation learnt from the spatiotemporal encoder, we use the Multidimensional Scaling (MDS) algorithm to reduce high-dimensional data to two dimensions while preserving inherent structural information and latent relationship (Wickelmaier,, 2003). This technique facilitates an intuitive visual representation of the data learnt by the deep learning models. For visualisation, we depict nonaccident and low-risk roads with green nodes, while blue nodes represent high-risk roads. The spatiotemporal embeddings, $\mathcal{Z}$, are illustrated in Figure 3.
(a) Lambeth
(b) Tower Hamlets
(c) Westminster
Figure 3: Low accident risk (green color) and high accident risk (blue color) spatio-temporal embedding $\mathcal{Z}$ MDS-visualisation on Lambeth, Tower Hamlets and Westminster with multi-step predict.
This visualisation reveals that roads associated with accidents are naturally clustered yet overlap with some low-risk areas, which introduces aleatoric uncertainties to the representation learning. Thus, traditional models may not be suitable in this case (Zhuang et al.,, 2022). Furthermore, this overlap can be attributed to spatial spillover effects, where adjacent roads exhibit similar embeddings despite their different risk levels. Although low- and high-risk roads with accident clusters are not distinct, high-risk roads within the Westminster data set are more discernible. Both the Lambeth and Tower Hamlets datasets show spatial aggregations that partially reflect the severity of accident risk, indicating that the requirements for an advanced model are to effectively incorporate accident risk severity in its embeddings and to be capable of capturing the complex long-tail distribution of risk within the accident data.
@@ 4.2 Experiment Setup
The parameters of STZITD-GNN are optimised with Adam optimiser (Kingma and Ba,, 2014), where the learning rate is set to 0.01, and weight decay is 0.01. In STZITD-GNN, $N_{epoch}=20$, the hidden unit is set to 42, GAT head is set to 3. The baseline parameters are optimised using Adam Kingma and Ba, (2014) with regularisation of $L_{2}$ and a dropout rate of 0.2. The GNNs in STZITD-GNNs and baselines are all two-layered. We also set the hidden units at 42. Similarly to previous work, we also employ the early stopping strategy with patience equal to 10 to avoid overfitting(Zhou et al., 2020a, ). STZITD-GNNs is implemented in Pytorch 1.9.0 with Python 3.8. All experiments were carried out on 1 NVIDIA GeForce RTX 3090, 24 GB.
@@ 4.3 Evaluation Metrics
We use eight metrics to evaluate the performance of our STZITD-GNNs from four perspectives, including the predictive regression effect of road accident risk scores, the predictive effect of uncertainty-aware distributional characteristics, the accuracy of zero accident road identification, and the accurate identification of accidents occurring on predicted high-risk roads.
4.3.1 Accident Accuracy Metrics
For the aspect of the point estimate, we compare the prediction of the numerical accident risk scores. For probabilistic distribution outputs like our model, we use the mean value of the distribution as the numerical output. For specific metrics, we use Mean Absolute Error (MAE), Mean Absolute Percentage error (MAPE), and Root Mean Squared Error (RMSE) to evaluate the predicted accident risk scores of each road. Specific definitions are summarized:
@@ 4.3.2 Uncertainty Quantification Metric
4.3.3 Zero Accident Identification Metrics
Given the problem of zero-value data sparsity, highlighted in Figure 2, it is essential to assess the ability of a model to accurately represent roads without accidents. We used the true zero rate (ZR) as a metric to quantify whether a model reflects the zero sparsity seen in the actual data. The formula is as follows:
@@ 4.3.3 Zero Accident Identification Metrics
4.3.4 Accuracy Hit Metric
To evaluate the effectiveness of our model in identifying roads with a higher likelihood of accident occurrence, we use the Accuracy at $a$ (Acc@$a$) metric. Specifically, we focus on the top 20% ($a=20\%$) of roads considered high-risk to verify if these areas indeed correspond to actual accidents. This alignment of predicted high-risk areas with real accident sites, as reported in Zhang and Cheng, (2020); Zhang et al., 2019b , allows Acc@$a$ to effectively measure the precision of our model in forecasting roads at increased risk of accidents. The formula for calculating Acc@$a$ is as follows:
$Acc@a=\frac{1}{p}\sum_{j=1}^{p}\frac{\sum_{i=1}^{N}\mathcal{I}(y_{ij}>0\cap\hat{y}_{ij}\geq\text{Top}(\hat{y}_{:j},a))}{\sum_{i=1}^{N}\mathcal{I}(y_{ij}>0)},$ (20)
The indicator function, $\mathcal{I}(y_{ij}>0\cap\hat{y}{ij}\geq\text{Top}(\hat{y}{:j},a))$, is rated as 1 when an accident (denoted by $y_{ij}>0$ for an occurrence on the $i^{th}$ road at the $j^{th}$ time slot) aligns with the top $a$ percent of predicted risks (where $\hat{y}{ij}$ meets or exceeds the threshold $\text{Top}(\hat{y}{:j},a)$), and 0 otherwise.
@@ 4.4 Baseline Models
In evaluating our proposed model, we carefully select baseline models that allow evaluation of the features of the proposed model: uncertainty quantification in multistep road-level accident risk prediction. Our selection criteria favour traditional point estimation methods and the most recent interval prediction models. This approach is in line with recent studies Zhuang et al., (2022). Specifically, our baselines include the mean-predict method (HA) and advanced spatial-temporal graph deep learning techniques such as STGCN and STGAT. Additionally, we consider parameterisation approaches such as STG-GNN, STNB-GNN, STTD-GNN, and STZINB-GNN. These probabilistic models have two or three parameter distributions, simpler than the four-parameter models we propose. The comparison models are:
  • HA(Liu and Guan, (2004)): The historical average serves as the minimum baseline. It is calculated by averaging road traffic accidents in the same time intervals from historical data to predict the future road traffic accident.
@@
  • STG-GNN(Wang et al., (2023)): Same as STZITD-GNN, we replace the statistical assumption of road traffic accidents with Gaussian distribution.
  • STNB-GNN(Zhuang et al., (2022)): Same as STZITD-GNN, we replace the statistical assumption of road traffic accidents with Negative Binomial Distribution.
@@
  • STZINB-GNN(Zhuang et al., (2022)): Unlike STZITD-GNN, we replace the statistical assumption of road traffic accidents with a zero-inflated negative binomial distribution.
@@ 4.5 Performance Comparison
We evaluated the performance of our proposed model, STZITD-GNN, against several baseline models from four different aspects: accuracy of point estimation, uncertainty quantification, non-risk road identification and accurate prediction of accident occurrence, as shown in Tables 4. The results are indicated by upward arrows, representing improvements over the second-best models. These enhancements confirm the significant impact of probabilistic assumptions on the accuracy and stability of our results, consistent with recent findings by Wang et al., (2023). The consistent performance of the STZITD-GNN model underscores the critical importance of selecting appropriate distributional assumptions that align with the characteristics of traffic accident data.
@@
For point estimation accuracy, the superior performance of the STZITD-GNN model across all regression error metrics highlights the effectiveness of the ZITD distribution. This distribution captures the long-tail dynamics of road traffic accident risk data with greater accuracy, making the mean value derived from the four parameters of the distribution decoder more representative of the actual risk scores. STZITD-GNN significantly improves the accuracy of the estimation of the accident risk score, demonstrating a 49. 06% increase in MAPE compared to the second-best model in the Lambeth case study, which has the most significant zero-inflation. Traditional deterministic models like HA, STGCN, and STGAT show varied performance, with HA consistently underperforming. GAT-based models, utilising predefined adjacency matrices that embed domain spatial knowledge, generally outperform GCN models by incorporating spatial proximity and contextual similarity, thereby better representing spatial dependencies. Among probabilistic models, those that assume Gaussian distributions perform poorly. They are significantly outperformed by deterministic models and show far worse performance 50% compared to the STZITD-GNN model, which uses skewed distribution assumptions. This underscores the importance of selecting appropriate distributional assumptions for probabilistic models.
@@
Second, the variation in uncertainty quantification across different models highlights the significant impact of distributional assumptions. The four-parameter distribution model (STZITD-GNN) shows substantial improvements, most notably in effectively capturing a narrow distribution while encompassing a wider range of actual accident risk values. In contrast, the worst-performing model, STG-GNNs, illustrates how improper distributional assumptions can degrade the quality of the estimation, introducing greater uncertainty into the results. The MPIW of this model is approximately 1.5 times that of the least uncertain three-parameter model, due to its constraints that enforce equal mean and variance in the distribution. Although ideally, outcome distributions would have a narrow interval, traditionally a wider MPIW suggests a higher PICP, indicating a trade-off in measuring uncertainty(Wang et al.,, 2023; Zhu et al.,, 2022). The STZINB-GNN and STTD-GNN models show that in scenarios characterised by extreme zero-inflation and heavier tails, a higher MPIW is associated with better PICP performance. This correlation suggests that broader prediction intervals, which encompass more observations, are beneficial in more extreme cases as they cover a wider range of observed values, despite reducing the ability to achieve narrower intervals. However, the STZITD-GNN model shows a promising improvement in MPIW, ranging from 47.30% to 55.07% in all case studies, highlighting its efficacy. In terms of PICP, it also shows a marginal increase of approximately 0.2% over the next best model, STTD-GNNs. This modest improvement still validates the impact of leveraging long-tailed and zero-inflated TD distributions, which contribute to a larger PICP and reduced variance in the predicted outcomes.
@@
In scenarios characterised by a high prevalence of non-accident roads, the STZITD-GNN model not only aims to precisely quantify accident risk values but also effectively differentiates non-accident roads. The integration of the sparsity parameter ($\pi$) significantly enhances the ability of the traditional TD assumption to address the prevalent zeros in the dataset, thereby reducing uncertainties associated with zero outcomes. The TD and ZINB distributions provide distinct parametric approaches to tackle the issue of zero-inflation. The TD distribution naturally integrates zero-inflation through its end-to-end structure, while the ZI parameter in the ZINB model requires an additional interpretative step like ZITD model does. For example, in the Tower Hamlets and Westminster cases, the ZI parameter significantly increases performance in zero-rate predictions compared to the NB distribution, showing improvements of 16.54% and 17.94%, respectively. In extreme zero-inflation scenarios, such as in Lambeth, both the STTD and STZINB models perform poorly, with a 15.09% discrepancy in ZR results compared to the STZITD-GNN model. Although the ZI paradigm is integrated into both the TD and the NB models, the ZITD model has a continuous parameter space, which can be advantageous for fitting data with a mix of zero inflation and continuous positive values. In contrast, traditional ZINB models operate under a discrete probabilistic assumption and struggle with complex data structures, particularly when the positive values are continuous rather than merely counts.
@@
In analyses that correlate high-risk roads with actual accidents, the STZITD-GNN model consistently demonstrates robust performance. In particular, it achieved a significant 76.59% precision in identifying actual accidents in the Lambeth case. Across all three distinct cases, the model maintains a robust and accurate identification rate, with a variance of only 7.61%. In contrast, models employing unsuitable distribution assumptions, such as STG-GNN, tend to significantly underpredict accident occurrences, often yielding the poorest results. This discrepancy underscores the importance of selecting the appropriate distribution model again.
@@
Non Accident Quantification ZR 0.5203 0.5305 0.5986 0.1992 0.5621 0.6838 0.6408 $\mathbf{0.7870}_{\uparrow 15.09\%}$
Accurate $Acc@20$ 0.4520 0.6113 0.6422 0.2666 0.4471 0.7123 0.6184 $\mathbf{0.7659}_{\uparrow 07.52\%}$
@@
Non Accident Quantification ZR 0.4998 0.4896 0.5234 0.1351 0.5158 0.6420 0.7526 $\mathbf{0.8876}_{\uparrow 17.94\%}$
Accurate $Acc@20$ $0.4752$ 0.5869 0.6950 0.2647 0.5022 0.6368 0.5827 $\mathbf{0.7224}_{\uparrow 03.94\%}$
@@
Non Accident Quantification ZR 0.5981 0.1431 0.2922 0.2071 0.5393 0.5079 0.6288 $\mathbf{0.7328}_{\uparrow 16.54\%}$
Accurate $Acc@20$ 0.4217 0.5020 0.4808 0.2933 0.4503 0.6075 0.5139 $\mathbf{0.6898}_{\uparrow 13.55\%}$
4.5.1 Visualization of Predicted Road-Level Accident
To illustrate the results, we have generated choropleth maps of road-level accidents across three distinct boroughs over a 14-day period. Considering the sparsity of accident data, which complicates the visualisation, we selected the three most risk-intensive periods of the data for display, as shown in Figure 4. Each borough reveals distinct accident patterns that correlate closely with the model’s predictions, offering a granular view of its predictive accuracy.
These visualizations highlight the general sparsity of traffic accident data, with most roads being non-risk (blue). Only a few roads exhibit higher accident risk levels (red), while others are associated with low-risk accidents. The STZITD-GNN model excels in identifying a range of risk levels, from zero-risk to high-risk. It performs particularly well in Westminster, accurately identifying a broader range of risk-prone roads, aided by the area’s lower zero inflation rate and denser graph structure. Furthermore, the model consistently and accurately maps the spatial connectivity of high-risk accident roads across all three boroughs, with risk areas typically clustering along major thoroughfares such as A11 Whitechapel Road and A13 in Tower Hamlets and A3 in Lambeth.
In the three distinct boroughs, the STZITD-GNN model’s ability to handle diverse spatial patterns and variations in daily accidents demonstrates its adaptability and consistent performance. However, the model occasionally misclassifies zero-risk roads as low-risk, particularly in Lambeth and Tower Hamlets. Although these errors generally fall within lower-risk categories and do not significantly impact the identification of higher-risk roads with actual accident occurrences, refining this aspect of the model could further improve its overall accuracy.
Figure 4: Predicted long-term (14-day) accidents on Lambeth, Tower Hamlets, and Westminster.

### v4

@@ 4.1 Traffic Crash Data Description
(b) The distribution of 2019-crash risk scores frequency in the three boroughs
Figure 2: Traffic crashes are depicted for the case studies.
The three high-crash frequency boroughs of Greater London, UK are illustrated in Figure 2. These boroughs each averaged 3 to 5 daily crashes throughout 2019, as detailed in Table 3. Despite these figures, a significant number of roads in these boroughs report no crashes, leading to a pronounced zero-inflation problem. This situation is compounded by the distribution of crash risk values (Figure 2), which shows a zero-inflated problem along with a pronounced long-tailed effect -where a few high-risk crash occurrences contrast with many low-risk ones. Such data characteristics underscore the suitability of our proposed model for this dataset, effectively capturing both the prevalent zero rates and the diverse spread of crash risk levels.
Table 3: The statistics of Westminster, Lambeth, and Tower Hamlets datasets in 2019.
Dataset  Roads/Nodes   Edges   Zero-Inflated
@@ 4.2 Experiment Setup
The train, validation and test data are all from 2019, as the ratio is 8:2:2. The parameters of STZITD-GNN are optimised with the Adam optimiser (Kingma and Ba,, 2014), where the learning rate is set to 0.01, and weight decay is 0.01. In STZITD-GNN, $N_{epoch}=20$, the hidden unit is set to 42, GAT head is set to 3. The baseline parameters are optimised using Adam Kingma and Ba, (2014) with regularisation of $L_{2}$ and a dropout rate of 0.2. The GNNs in STZITD-GNNs and baselines are all two-layered. We also set the hidden units at 42. Similarly to previous work, we also employ the early stopping strategy with patience equal to 10 to avoid overfitting (Zhou et al., 2020a, ). STZITD-GNNs is implemented in Pytorch 1.9.0 with Python 3.8. All experiments were carried out on 1 NVIDIA GeForce RTX 3090, 24 GB.
@@ 4.3 Evaluation Metrics
We use eight metrics to evaluate the performance of our STZITD-GNNs from four perspectives, including the predictive regression effect of road crash risk scores, the predictive effect of uncertainty-aware distributional characteristics, the accuracy of zero crash road identification, and the accurate identification of crashes occurring on predicted high-risk roads.
4.3.1 Crash Accuracy Metrics
For the aspect of the point estimate, we compare the prediction of the numerical crash risk scores. For probabilistic distribution outputs like our model, we use the mean value of the distribution as the numerical output. For specific metrics, we use Mean Absolute Error (MAE), Mean Absolute Percentage error (MAPE), and Root Mean Squared Error (RMSE) to evaluate the predicted crash risk scores of each road. Specific definitions are summarized:
@@ 4.3.2 Uncertainty Quantification Metric
4.3.3 Zero crash Identification Metrics
Given the problem of zero-value data sparsity, highlighted in Figure 2, it is essential to assess the ability of a model to accurately represent roads without crashes. We used the true zero rate (ZR) as a metric to quantify whether a model reflects the zero sparsity seen in the actual data. The formula is as follows:
@@ 4.3.3 Zero crash Identification Metrics
4.3.4 Accuracy Hit Rate Metric on Crash Occurence
To evaluate the effectiveness of our model in identifying roads with a higher likelihood of crash occurrence, we use the accuracy hit rate in $a$ ($AccHR@20$) metric. Specifically, we focus on the top 20% ($a=20\%$) of roads considered high-risk to verify if these areas actually correspond to actual crash occurrence. This alignment of predicted high-risk areas with real crash sites, as reported in Zhang and Cheng, (2020); Zhang et al., 2019b , allows Acc@$a$ to effectively measure the precision of our model in forecasting roads at increased risk of crashes. The formula for calculating Acc@$a$ is as follows:
$AccHR@a=\frac{1}{p}\sum_{j=1}^{p}\frac{\sum_{i=1}^{N}\mathcal{I}(y_{ij}>0\cap\hat{y}_{ij}\geq\text{Top}(\hat{y}_{:j},a))}{\sum_{i=1}^{N}\mathcal{I}(y_{ij}>0)},$ (16)
The indicator function, $\mathcal{I}(y_{ij}>0\cap\hat{y}{ij}\geq\text{Top}(\hat{y}{:j},a))$, is rated as 1 when a crash (denoted as $y_{ij}>0$ for a crash occurrence on the $i^{th}$ road in the $j^{th}$ time interval) aligns with the top $a$ percent of predicted risks (where $\hat{y}{ij}$ meets or exceeds the threshold $\text{Top}(\hat{y}{:j},a)$), and 0 otherwise.
@@ 4.4 Baseline Models
In evaluating our proposed model, we carefully select baseline models that allow evaluation of the features of the proposed model: uncertainty quantification in multistep road-level crash risk prediction. Our selection criteria favour traditional point estimation methods and the most recent interval prediction models. This approach is in line with recent studies Zhuang et al., (2022). Specifically, our baselines include the mean-predict method (HA) and advanced spatial-temporal graph deep learning techniques such as STGCN and STGAT. Additionally, we consider parameterisation approaches such as STG-GNN, STNB-GNN, STTD-GNN, and STZINB-GNN. These probabilistic models have two or three parameter distributions, simpler than the four-parameter models we propose. The comparison models are:
  • HA(Liu and Guan, (2004)): The historical average serves as the minimum baseline. It is calculated by averaging road traffic crashes in the same time intervals from historical data to predict the future road traffic crash.
@@
  • STG-GNN(Wang et al., (2023)): Same as STZITD-GNN, we replace the statistical assumption of road traffic crashes with Gaussian distribution.
  • STNB-GNN(Zhuang et al., (2022)): Same as STZITD-GNN, we replace the statistical assumption of road traffic crashes with Negative Binomial Distribution.
@@
  • STZINB-GNN(Zhuang et al., (2022)): Unlike STZITD-GNN, we replace the statistical assumption of road traffic crashes with a zero-inflated negative binomial distribution.
@@ 4.5 Performance Comparison
We evaluated the performance of our proposed model, STZITD-GNN, against several baseline models from four different aspects: accuracy of point estimation, uncertainty quantification, non-risk road identification and accurate prediction of crash occurrence, as shown in Tables 4. The results are indicated by upward arrows, representing improvements over the second-best models. These enhancements confirm the significant impact of probabilistic assumptions on the accuracy and stability of our results, consistent with recent findings by Wang et al., (2023). The consistent performance of the STZITD-GNN model underscores the critical importance of selecting appropriate distributional assumptions that align with the characteristics of traffic crash data.
@@
For point estimation accuracy, the superior performance of the STZITD-GNN model across all regression error metrics highlights the effectiveness of the ZITD distribution. This distribution captures the long-tail dynamics of road traffic crash risk data with greater accuracy, making the mean value derived from the four parameters of the distribution decoder more representative of the actual risk scores. STZITD-GNN significantly improves the accuracy of the estimation of the crash risk score, demonstrating a 49.06% increase in MAPE compared to the second-best model in the Lambeth case study, which has the most significant zero-inflation. Traditional deterministic models like HA, STGCN, and STGAT show varied performance, with HA consistently underperforming. GAT-based models, utilising predefined adjacency matrices that embed domain spatial knowledge, generally outperform GCN models by incorporating spatial proximity and contextual similarity, thereby better representing spatial dependencies. Among probabilistic models, those that assume Gaussian distributions perform poorly. They are significantly outperformed by deterministic models and show far worse performance 50% compared to the STZITD-GNN model, which uses skewed distribution assumptions. This underscores the importance of selecting appropriate distributional assumptions for probabilistic models.
@@
Second, the variation in uncertainty quantification across different models highlights the significant impact of distributional assumptions. The four-parameter distribution model (STZITD-GNN) shows substantial improvements, most notably in effectively capturing a narrow distribution while encompassing a wider range of actual crash risk values. In contrast, the worst-performing model, STG-GNNs, illustrates how improper distributional assumptions can degrade the quality of the estimation, introducing greater uncertainty into the results. The MPIW of this model is approximately 1.5 times that of the least uncertain three-parameter model, due to its constraints that enforce equal mean and variance in the distribution. Although ideally, outcome distributions would have a narrow interval, traditionally a wider MPIW suggests a higher PICP, indicating a trade-off in measuring uncertainty (Wang et al.,, 2023; Zhu et al.,, 2022; Zhuang et al.,, 2022). The STZINB-GNN and STTD-GNN models show that in scenarios characterised by extreme zero-inflation and heavier tails, a higher MPIW is associated with better PICP performance. This correlation suggests that broader prediction intervals, which encompass more observations, are beneficial in more extreme cases as they cover a wider range of observed values, despite reducing the ability to achieve narrower intervals. However, the STZITD-GNN model shows a promising improvement in MPIW, ranging from 47.30% to 55.07% in all case studies, highlighting its efficacy. In terms of PICP, it also shows a marginal increase of approximately 0.2% over the next best model, STTD-GNNs. This modest improvement still validates the impact of leveraging long-tailed and zero-inflated TD distributions, which contribute to a larger PICP and reduced variance in the predicted outcomes.
@@
In scenarios characterised by a high prevalence of non-crash roads, the STZITD-GNN model not only aims to precisely quantify crash risk values but also effectively differentiates non-crash roads. The integration of the sparsity parameter ($\pi$) significantly enhances the ability of the traditional TD assumption to address the prevalent zeros in the dataset, thereby reducing uncertainties associated with zero outcomes. The TD and ZINB distributions provide distinct parametric approaches to tackle the issue of zero-inflation. The TD distribution naturally integrates zero-inflation through its end-to-end structure, while the ZI parameter in the ZINB model requires an additional interpretative step like ZITD model does. For example, in the Tower Hamlets and Westminster cases, the ZI parameter significantly increases performance in zero-rate predictions compared to the NB distribution, showing improvements of 16.54% and 17.94%, respectively. In extreme zero-inflation scenarios, such as in Lambeth, both the STTD and STZINB models perform poorly, with a 15.09% discrepancy in ZR results compared to the STZITD-GNN model. Although the ZI paradigm is integrated into both the TD and the NB models, the ZITD model has a continuous parameter space, which can be advantageous for fitting data with a mix of zero inflation and continuous positive values. In contrast, traditional ZINB models operate under a discrete probabilistic assumption and struggle with complex data structures, particularly when the positive values are continuous rather than merely counts.
@@
In analyses that correlate high-risk roads with actual crashes, the STZITD-GNN model consistently demonstrates robust performance. In particular, it achieved a significant 76.59% precision in identifying actual crashes in the Lambeth case. Across all three distinct cases, the model maintains a robust and accurate identification rate, with a variance of only 7.61%. In contrast, models employing unsuitable distribution assumptions, such as STG-GNN, tend to significantly underpredict crash occurrences, often yielding the poorest results. This discrepancy underscores the importance of selecting the appropriate distribution model again.
@@
Non crash Quantification ZR 0.5203 0.5305 0.5986 0.1992 0.5621 0.6838 0.6408 $\mathbf{0.7870}_{\uparrow 15.09\%}$
Hit Rate $AccHR@20$ 0.4520 0.6113 0.6422 0.2666 0.4471 0.7123 0.6184 $\mathbf{0.7659}_{\uparrow 07.52\%}$
@@
Non crash Quantification ZR 0.4998 0.4896 0.5234 0.1351 0.5158 0.6420 0.7526 $\mathbf{0.8876}_{\uparrow 17.94\%}$
Hit Rate $AccHR@20$ $0.4752$ 0.5869 0.6950 0.2647 0.5022 0.6368 0.5827 $\mathbf{0.7224}_{\uparrow 03.94\%}$
@@
Non crash Quantification ZR 0.5981 0.1431 0.2922 0.2071 0.5393 0.5079 0.6288 $\mathbf{0.7328}_{\uparrow 16.54\%}$
Hit Rate $AccHR@20$ 0.4217 0.5020 0.4808 0.2933 0.4503 0.6075 0.5139 $\mathbf{0.6898}_{\uparrow 13.55\%}$
4.5.1 Visualization of Predicted Road-Level crash
To illustrate the results, we have generated Choropleth maps of road-level crashes across three distinct boroughs over a 14-day period. Considering the sparsity of crash data, which complicates the visualisation, we selected the three most risk-intensive periods of the data for display, as shown in Figure 3. Each borough reveals distinct crash patterns that correlate closely with the model’s predictions, offering a granular view of its predictive accuracy.
These visualizations highlight the general sparsity of traffic crash data, with most roads being non-risk (blue). Only a few roads exhibit higher crash risk levels (red), while others are associated with low-risk crashes. The STZITD-GNN model excels in identifying a range of risk levels, from zero-risk to high-risk. It performs particularly well in Westminster, accurately identifying a broader range of risk-prone roads, aided by the area’s lower zero inflation rate and denser graph structure. Furthermore, the model consistently and accurately maps the spatial connectivity of high-risk crash roads across all three boroughs, with risk areas typically clustering along major thoroughfares such as A11 Whitechapel Road and A13 in Tower Hamlets and A3 in Lambeth.
In the three distinct boroughs, the STZITD-GNN model’s ability to handle diverse spatial patterns and variations in daily crashes demonstrates its adaptability and consistent performance. However, the model occasionally misclassifies zero-risk roads as low-risk, particularly in Lambeth and Tower Hamlets. Although these errors generally fall within lower-risk categories and do not significantly impact the identification of higher-risk roads with actual crash occurrences, refining this aspect of the model could further improve its overall accuracy.
Figure 3: Real v.s. Predicted Crash Risks on Lambeth, Tower Hamlets, and Westminster.